Solve using a system of equations (show the two equations you used then show the solutions):
Tickets for a concert sold for $11 and $8. If 600 tickets were sold for one evening and the receipts are $5550, how many of each kind of ticket were sold?

Answer by checkley77(12844)

X+Y=600 OR X=600-Y
11X+8Y=5550
11(600-Y)+8Y=5550
6600-11Y+8Y=5550
-3Y=5550-6600
-3Y=-1050
Y=-1050/-3
Y=350 $8 TICKETS WERE SOLD.
X+350=600
X=600-350
X=250 $11 TICKETS WERE SOLD.
PROOF:
11*250+8*350=5550
2750+2800=5550
5550=5550